OpenStudy (anonymous):

5/2+square root on 15???

OpenStudy (anonymous):

\[\frac{5}{2+\sqrt{15}}\] multiply top and bottom by \[2-\sqrt{15}\]

OpenStudy (anonymous):

denominator will be \[4-15=-11\] numerator will be whatever you get

OpenStudy (anonymous):

so wats the numerator?

OpenStudy (anonymous):

I'm confused what the question is.